The period of a certain pendulum is 2.0 seconds and the mass of the pendulum Bob is 50g. The Bob is pulled aside through a horizontal distance of 8cm and then released. Find the displacement and kinetic energy of the Bob 0.7seconds after release


1
Expert's answer
2021-02-10T07:16:57-0500

Answer

Angular frequency

"w=\frac{2\pi }{T}=3.14" rad/sec

t=0.7sec

x0=0.08m

Amplitude is given

"X=x_0sin wt"

"X=0.08\times sin(3.14\times0.7)"

X=0.064m

Now velocity

"V=X\times w=0.2" m/s

So kinetic energy

"K=\frac{mV^2}{2}=4\times10^{-5}" J